OpenStudy (anonymous):

Medal Awarded (: What effect does a value of h = 5 have on the graph of a radical function?

OpenStudy (anonymous):

a. Vertical Shift of 5 units up. b. Vertical Shift of 5 units to the left. c. Vertical Shift of 5 units to the right. d. Vertical Shift of 5 units down.

OpenStudy (anonymous):

General form\[y=a \sqrt{b(x-h)}+k\]

OpenStudy (anonymous):

Check this out

OpenStudy (anonymous):

(y-h) = m(x-k) + c By substituting y for y - 5 (as h=5), it shifts the graph 5 steps downwards (as every y value is subtracted by 5)

OpenStudy (anonymous):

im confused...

OpenStudy (anonymous):

Open the attached pic